Reports 1-1 of 1 Clear search Modify search
Environmental Monitoring (Environmental Monitoring)
Tringali, Fiori, Spinicelli, Zaza - 0:40 Wednesday 12 November 2025 (68143) Print this report
Comment to Laser Lab acoustic injection: setup and tests (67772)

The purpose of this shift was to test a set of values sent to the DAC channel ENV_Noise_CEB_EE_room.
Only two values were tested due to some issues with the lock stability.

Each injection lasted 180 seconds and was performed via the METATRON node.

Some comments are reported below:

  • The ITF unlock was not caused by the acoustic injections (see Figure 1).

  • Figure 2 shows several monitoring signals from the BPC (BsX_*), the microphones, and their coherence with Hrec, while Figure 3 displays the same signals with a zoom around 800 Hz. At this frequency, Hrec exhibits a coherent structure with some of the BPC signals, suggesting a possible coupling path.

  • Figure 4 shows the DAC noise levels.

  • Figures 5 and 6 show the acoustic spectra during the injections.

  • Figures 7 and 8 show the vibration of the benches during the injections.

  • Figure 9 shows the structures excited in Hrec in the frequency range 100–1000 Hz.

     

Value Time (UTC)
0.01 17:48:20
0.02 17:54:48
---  ITF unlock at 17:59

 

Images attached to this comment
Search Help
×

Warning

Error

The present report has been modified outside this window. Please check for its integrity in the main page.

Refreshing this page will move this report into drafts.

×